OpenStudy (anonymous):

0.5b + 4 = 2(b + 2) (1 point)

OpenStudy (anonymous):

Yes that seems right. Here's my work. \[\prod_{\eta \pi \chi =\subseteq \eta \chi \cot \sqrt{\sqrt[?]{?}}}^{?}\]

OpenStudy (anonymous):

Everyone look out for Lala4jc, she is trolling.

OpenStudy (anonymous):

I know!

OpenStudy (anonymous):

im not!

OpenStudy (anonymous):

now you're lying

OpenStudy (anonymous):

i really need help on this problemm

OpenStudy (anonymous):

b = 0 is the answer

OpenStudy (anonymous):

Well you have come to the right place

OpenStudy (anonymous):

That's called cheating, do your own work.

OpenStudy (anonymous):

0.5b + 4 = 2(b + 2) please im stuck

OpenStudy (anonymous):

Maybe you should have studied last night.

OpenStudy (anonymous):

well obviously it's 0.5b + 4 = 2b + 4

OpenStudy (anonymous):

LL4jc is right

OpenStudy (anonymous):

Ashleylynn the answer is b= 0

OpenStudy (anonymous):

which gives you zero nimrod